Can I get help to ((find p if (5,p) lies on the straight line with equation 3x+4y-5=0 ))
Answer by stanbon(75887)

Can I get help to ((find p if (5,p) lies on the straight line with equation 3x+4y-5=0 ))
--------
Substitute::
3*5 + 4p - 5 = 0
------
4p = -10
------
p = -5/2
------------
Cheers,
Stan H.
------------